# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

### v2

@@ IV.2 The infalling spherical accretion
Figure 10: The specific intensity $I_{\text{obs}}$ of a infalling spherical accretion as observed by a distant observer.
We next investigate the case of infalling spherical accretion around Schwarzschild-Hernquist black holes. The general formula for the observed intensity, Eq.(33), remains valid for infalling accretion. However, the redshift factor $\mathcal{R}$ must be modified to account for the velocity of the infalling matter. For infalling accretion, the redshift factor is given by Li and He (2021b, a); Fathi and Cruz (2023); Bambi (2013)
@@
Figure 11: The images of Schwarzschild-Hernquist black holes’ shaodow with infalling spherical accretion as observed by a distant observer.
@@
The observed intensity profile $I_{\text{obs}}$ and the corresponding images for infalling accretion, computed using Eq.(43), are shown in Figs.10 and 11, respectively.
As shown in Fig.10, and similar to the static case, the observed intensity $I_{\text{obs}}$ for infalling accretion peaks near the critical impact parameter $b_{p}$ for all Hernquist DM parameter. The intensity increases for $b<b_{p}$, rising sharply just before reaching the peak at $b\sim b_{p}$. For $b<b_{p}$, the intensity gradually decreases with increasing $b$, consistent with the behavior in the static model. The observed intensity at $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is substantially brighter than at $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, as shown in Figs.8 and 10. A direct comparison between the images in Fig.9 and Fig.11 reveals that the central region is significantly darker for infalling accretion. When the parameters are changed from $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ to $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, the peak intensity decreases by $18.36\%$ for static accretion and by $50.50\%$ for infalling accretion. As a direct result of the Doppler effect, the apparent intensity of radiation emitted by materials falling radially inward is reduced due to redshifting. Furthermore, in both the static and infalling models, the radius of the photon ring increases with increasing Hernquist DM parameters, consistent with the findings in Tab.1.
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadows of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. Employing ray-tracing techniques and numerical simulations, we have quantitatively analyzed three thin-disk accretion models (26)-(32), and two spherical accretion models Eqs.(37) and (43), revealing that the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ have a significant impact on the photon sphere radius and the observed shadow properties. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.6-11. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In extreme circumstances, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 9 and 11, the peak observed intensity is significantly diminished, giving the overall image of being substantially dimmer. The gravitational influence of the Hernquist DM halo is responsible for these alterations, modifying the effective potential around the black hole and the photon trajectory. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
@@
Importantly, our researches offer significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, offers a powerful diagnostic for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides an unambiguous prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, a future measurement of an anomalously large shadow could indicate the presence of a Hernquist DM halo, and our model could be used to interpret such a result. Conversely, the absence of significant shadow enlargement would place stringent constraints on the parameter space of viable Hernquist DM models.
@@
Looking forward, our results highlight the potential of next-generation very-long-baseline interferometry (VLBI) Hobiger et al. (2009); Schuh and Böhm (2012); Niell et al. (2018); Nothnagel et al. (2018) arrays to refine DM models through precise shadow imaging. Furthermore, they reveal a novel, observable coupling mechanism between DM and black hole shadow phenomenology. Future work incorporating black hole spin like the Kerr metric Ferrer et al. (2017); Hou et al. (2018b) and more realistic, dynamic accretion flows will be essential to fully bridge the gap between theoretical predictions and observational evidence. Our results strongly support the use of black hole shadows as unique probes that can reveal the distribution and properties of DM halos, as well as elucidate new aspects of gravitational physics in strong-field regimes. This approach promises to illuminate the nature of DM in galactic centers in unprecedented ways.

### v3

@@ IV.2 The infalling spherical accretion
Figure 11: The specific intensity $I_{\text{obs}}$ of an infalling spherical accretion as observed by a distant observer. The photon ring is represented by the sharp peak, similar to the static case,. The intensity is suppressed due to the Doppler de-boosting effect of the infalling matter. As the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ increase, the peak intensity decreases significantly, while the position of the peak shifts to larger impact parameters $b$.
We next investigate the case of infalling spherical accretion around Schwarzschild-Hernquist black holes. The general formula for the observed intensity, Eq.(39), remains valid for infalling accretion. However, the redshift factor $\mathcal{R}$ must be modified to account for the velocity of the infalling matter. For infalling accretion, the redshift factor is given by Li and He (2021a, b); Fathi and Cruz (2023); Bambi (2013)
@@
Figure 12: The images of Schwarzschild-Hernquist black holes’ shadow with infalling spherical accretion as observed by a distant observer. To enable direct quantitative comparison, all panels utilize a unified intensity color bar. The bright photon ring, which expands geometrically with increasing Hernquist DM parameters, outlines the shadow. Different from the static case, the central region exhibits pronounced darkness due to the Doppler de-boosting effect of the infalling matter.
@@
The observed intensity profile $I_{\text{obs}}$ and the corresponding images for infalling accretion, computed using Eq.(49), are shown in Figs.11 and 12, respectively.
As shown in Fig.11, and similar to the static case, the observed intensity $I_{\text{obs}}$ for infalling accretion peaks near the critical impact parameter $b_{p}$ for all Hernquist DM parameter. The intensity increases for $b<b_{p}$, rising sharply just before reaching the peak at $b\sim b_{p}$. For $b>b_{p}$, the intensity gradually decreases with increasing $b$, consistent with the behavior in the static model. The observed intensity at $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is substantially brighter than at $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, as shown in Figs.9 and 11. A direct comparison between the images in Fig.10 and Fig.12 reveals that the central region is significantly darker for infalling accretion. When the parameters are changed from $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ to $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, the peak intensity decreases by $18.36\%$ for static accretion and by $50.50\%$ for infalling accretion. This suppression is a kinematic radiative effect caused by Doppler de-boosting, where the apparent intensity of radiation emitted by materials falling radially inward is reduced due to redshifting. Contrary to these radiative fluctuations, which are model-dependent, the geometric signature is robust. In both the static and infalling models, the radius of the photon ring increases with increasing Hernquist DM parameters, consistent with the findings in Tab.1.
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadows of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.7-12. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In the limiting case, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 10 and 12, the peak observed intensity is significantly diminished, resulting in a substantially dimmer overall image. The geometric expansion of the photon trajectory is controlled by the Hernquist DM halo’s gravitational impact, although radiative transfer processes like gravitational redshift and Doppler de-boosting are the main causes of the intensity suppression. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
The parameter space explored in this study requires particular attention to its physical significance, especially the core density $\rho_{c}M^{2}\in[0.4,0.8]$ and core radius $r_{s}/M\in[0.2,0.4]$. To ensure that spacetime near the photon sphere is not dominated by halo contribution, we could estimate the Herquist DM mass at the typical radius $r=5M$. Employing the Hernquist DM mass (2), we find that the ratio of DM mass to BH mass, $(M_{\text{H}}/M_{\text{BH}})|_{r=5M}$, ranges from a negligible value $\sim 1.9\%$ for $r_{s}=0.2M$ and $\rho_{c}M^{2}=0.4$ to a significant yet sub-dominant $\sim 27.6\%$ for $r_{s}=0.4M$ and $\rho_{c}M^{2}=0.8$. This confirms that the geometry of the spacetime remains dominated by the central BH across the whole parameter space. We also compare our chosen parameters with Jha’s suggested limitations with respect to astronomical observations Jha (2025). The parameters used in this study for the supermassive BH $\text{M87}^{*}$ fall within the $2\sigma$ upper limits obtained from EHT data. The same parameters, however, far surpass the more stringent limitations for $\text{SgrA}^{*}$. Thus, it is crucial to explicitly classify this model is a phenomenological toy model that investigates theoretical upper bounds. However, the analysis’s findings still have definite ramifications for astronomical observations:
  • Exclusion Region: The expected shadow deviation approaches $30\%$ in high-density areas. The probability of such high-density Hernquist DM regions in the central region of the Milky Way is essentially ruled out by this large signal, which is much larger than current observational errors of $\sim 10\%$.
  • Feasibility Region: In contrast, the deviation becomes minor at $\sim 1.89\%$, falling below observational error limits, when parameters adopt lower values. This suggests that the lower-density Hernquist DM distributions are still viable astrophysical candidates for explaining observed anomalies under the limitations of present technology.
Future facilities like ngEHT will improve precision. The implications of this work might place more stringent upper constraints on Hernquist DM density around the BH if no deviation is found with better accuracy. On the other hand, the low-density parameter space shown here may offer a convincing physical explanation for non-Kerr features if tiny deviations are found.
@@
Importantly, our research offers significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, serves as a valuable theoretical framework for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides a theoretical prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, when an anomalously large shadow detected in the future indicate the presence of a Hernquist DM halo, our current analysis suggests that the absence of significant shadow enlargement already places stringent constraints on the parameter space of viable Hernquist DM models.
@@
Looking forward, our results highlight the potential of next-generation very-long-baseline interferometry (VLBI) Hobiger et al. (2009); Schuh and Böhm (2012); Niell et al. (2018); Nothnagel et al. (2018) arrays to refine DM models through precise shadow imaging. Furthermore, they demonstrate a distinct, observable correlation between DM parameters. Future work incorporating black hole spin in Kerr-like spacetimes Ferrer et al. (2017); Hou et al. (2018a) and more realistic, dynamic accretion flows will be essential to fully bridge the gap between theoretical predictions and observational evidence. Our results strongly support the use of black hole shadows as unique probes that can reveal the distribution and properties of DM halos, as well as elucidate new aspects of gravitational physics in strong-field regimes. This approach promises to shed new light on the nature of DM in galactic centers with increasing precision.